OpenStudy (anonymous):

Simplify: 3x^2-5x+2 ----------- x^2-3x+2 Work PLEASE! :D

OpenStudy (mertsj):

Did you try factoring and canceling out any common factors?

OpenStudy (anonymous):

yes. but i want to make sure i did it right.

OpenStudy (mertsj):

\[\frac{3x^2-5x+2}{x^2-3x+2}=\frac{(3x-2)(x-1)}{(x-2)(x-1)}=\frac{3x-2}{x-2}\]